# DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation

### v2

@@ F.3 Step 2: Claim Verification
Appendix G LLM-based Information Verification Implementation
G.1 Overview
@@
G.2 Claim Extraction and Classification
@@
The Information Verification Module automates the procedure of human evaluators. After tokenizing the report into sentences, the LLM evaluates the "claim likelihood" of each sentence and predicts A–F types based on the 6-Type definitions defined in Table 13. For Type B and C claims, evidence sentences are subsequently searched using Back-tracking.
@@
To mitigate the “Lost-in-the-Middle” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences belonging to the current batch ($s_{i},\dots,s_{i+B-1}$).
Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor. User: # Full Report Context {Full_Report_Text} # Target Sentences to Extract Claims From L10.S1: … L10.S20: … Extract claims only from the target sentences above. Use the full report context for coreference resolution.
@@
This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
Claim Extraction Evaluation Setup
To rigorously evaluate the claim extraction performance, we employed an LLM-based Judge (GPT-5). Standard lexical metrics (e.g., ROUGE Lin (2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning. We defined two key metrics: (1) Semantic Recall: Measures whether each ground truth claim is semantically covered by the set of extracted claims. The LLM Judge compares the ground truth claim against all candidate claims extracted from the same source sentence and determines if the core information is present. (2) Classification F1: Measures whether the LLM correctly classified the claim type for the extracted claims. The implementation code and LLM prompts used for the evaluation are included in the supplementary material.
Comparison with Sliding Window Baseline.
To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window” baseline on the subset of correctly classified B/C claims ($N=131$). The sliding window method collects all citations within a window of size $k$ centered on the claim. As shown in Table 15, the proposed LLM method achieves the highest Jaccard Index (0.7070) and Precision (0.7109), outperforming the sliding window baselines ($k=5,10,15$). While increasing the window size ($k$) improves Recall (up to 0.93), it significantly degrades Precision and Jaccard due to the inclusion of irrelevant citations. This result demonstrates that the LLM’s “Evidence Position” prediction provides a precise pointer to the supporting evidence, which is crucial for efficient verification.
@@
G.3 Claim Verification
Semantic Back-tracking for Citation Recovery
@@ G.3 Claim Verification
In this study, the BM25 Robertson et al. (2009) and OpenAI’s text-embedding-3-large OpenAI was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
@@
The Information Verification Module’s automatic verifier is prompt-engineered to determine whether the given context supports the claim, identical to the human protocol above. The LLM infers whether the claim’s core content and numerical information match the source to make a final judgment.
@@
To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 Robertson et al. (2009), OpenAI’s text-embedding-3-large OpenAI ), and Context Size (Top-K=2, 4). Table 16 summarizes the results on both the Original and Adversarial (Augmented) datasets. We observed that increasing the retrieved context size from Top-K=2 to 4 improved accuracy on the Original dataset (e.g., 77.0% $\rightarrow$ 79.3% for OpenAI, Low Effort), but increased the cost by approximately 35% (\$0.95 $\rightarrow$ \$1.28). The Low Reasoning Effort setting proved to be highly cost-effective, achieving comparable or superior performance to Medium/High effort while costing significantly less. Notably, the model demonstrated high robustness on the Adversarial dataset, maintaining high accuracy (>>88%) across most configurations, suggesting that it effectively distinguishes unsupported claims even under perturbations. Consequently, prioritizing feasibility of large-scale verification (cost/throughput) while retaining strong robustness, we selected the configuration Batch 20, Low Effort, Top-K=2, and OpenAI Embedding. This setup offers a minimal cost of \$0.95 per 1k claims while maintaining a strong accuracy of 77.0% (Original) and 88.5% (Adversarial), making it the most balanced choice for our resource-constrained high-volume verification pipeline.
@@
$\Rightarrow$ Final Result: Supported
G.4 Evaluation Metrics
The Information Verification Module calculates the following detailed metrics to evaluate Integrity and Sufficiency.
@@ G.4 Evaluation Metrics
  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index Rhoades (1993).
@@ 2. Aggregation
Appendix H Baseline Model Details
We use the following backbone model families in our experiments: Qwen3-235B, Gemini 2.5, Claude Opus 4.5, and GPT-5. For readability, we refer to the GPT family as GPT-5 in the paper, while the actual backbone used in our runs was GPT-5.2. Since Gemini 2.5 Pro includes reasoning by default, we use Gemini 2.5 Flash for the fast (non-reasoning) setting, and Gemini 2.5 Pro for the other settings. For think, we use each service’s default reasoning budget. For think+search, we do not build a custom retrieval pipeline; instead, we use the built-in web search system provided by each service. We exclude Gemini think+search because citation information is not provided in its outputs. For WebThinker, we use Qwen3-235B as an auxiliary model. For OpenAI Deep Research, we collect reports generated from the service environment as of August 2025.
Appendix I Human-Correlation Experiment Setup
We measure alignment between LLM judges and expert human judgments on 45 reports. We consider five domains and sample three tasks from each, yielding 15 tasks in total. For each task, we randomly sample three reports from those produced by five Deep Research systems—OpenAI Deep Research, Gemini 2.5 Pro Deep Research, Claude Opus 4.1 Deep Research, WebThinker, and Qwen3-235B Deep Research—resulting in $15\times 3=45$ reports overall. WebThinker Li et al. (2025b) uses Qwen3-235B as an auxiliary model.
@@ Appendix I Human-Correlation Experiment Setup
To measure alignment with expert human judgments, we report Pearson correlation ($r$), Spearman rank correlation ($\rho$), and pairwise agreement (PA). Each task corresponds to a single query and contains three reports. For each task, we compute Pearson $r$ and Spearman $\rho$ between human and model scores across the three reports, and report the mean over the 15 tasks. Following DeepBench Du et al. (2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs. We compute PA per task and report the mean across tasks. To match the human 1–5 scale, we rescale model scores to 1–5 by dividing 1–10 scores by 2. For PA, a pair is counted as an agreement if the model and human judgments induce the same relation (>>, <<, or ==) for that pair. Tasks with undefined correlations (e.g., zero variance) are excluded from the corresponding averages.
@@
Appendix J Ethical Considerations and Potential Risks
DEER is designed as a research evaluation benchmark rather than a deployment or decision-making system, and its scores should be interpreted accordingly. Because the framework partially relies on LLM-based judges, evaluation outcomes may reflect systematic biases or limitations of the underlying models; however, this risk is mitigated through expert-authored rubrics, task-specific expert guidance, and validation against human expert judgments. Unless otherwise specified, reported results correspond to single-run evaluations under fixed configurations, rather than aggregates over multiple random seeds. DEER evaluates the quality and evidence alignment of generated reports, not the real-world correctness, safety, or deployability of the underlying claims or recommendations, and thus should not be used as an endorsement for practical application in high-stakes domains. In addition, while the information verification module checks both explicit and implicit claims against external sources, its reliability is inherently bounded by the quality, availability, and temporal stability of retrieved web evidence. Overall, DEER is intended for comparative research analysis and diagnostic evaluation, and any use beyond this scope should be accompanied by appropriate human oversight and domain-specific safeguards.
Appendix K Information About Use of AI Assistants
AI assistants were used in a limited and supportive manner during the preparation of this work. Specifically, they were employed to assist with coding-related tasks (e.g., drafting boilerplate scripts, refactoring, and debugging) and with writing-related support, such as grammar checking, stylistic refinement, and clarity improvement. Importantly, all conceptual contributions, experimental design, analysis, interpretation of results, and final manuscript were carried out by the authors.

### v3

@@ E.3 Step 2: Claim Verification
Appendix F LLM-based Information Verification Implementation
F.1 Overview
@@
F.2 Claim Extraction and Classification
@@
The Information Verification Module is designed to automate the human verification protocol described above. The LLM analyzes the report sentence by sentence to extract and classify claims, and, if necessary, retrieves external documents to verify their factual accuracy. In this process, algorithms such as Batch Extraction, Back-tracking, and Relevant Context Filtering were applied to achieve both cost efficiency and accuracy.
@@
To mitigate the “Lost-in-the-Middle (Liu et al., 2023a)” phenomenon that occurs when processing long contexts and to increase cost efficiency, this study introduced a Batch Extraction strategy. After dividing the entire report $D$ into sentence units $S=\{s_{1},s_{2},\dots,s_{m}\}$, they are processed in batches of a fixed size $B$ (in this study, $B=20$). Each batch of processing provides the full report context ($D$) at the beginning of the prompt, but instructs the model to extract claims only for the sentences in the current batch ($s_{i},\dots,s_{i+B-1}$). Figure 7 shows the simplified prompt structure used for batch extraction, along with an example JSON output. This method induces the model to focus on local sentences while remaining aware of the entire context, achieving human-level claim-extraction performance.
Simplified Batch Extraction Prompt Structure System: You are an expert fact-checker and claim extractor.
User:
# Full Report Context
{Full_Report_Text}
# Target Sentences to Extract Claims From
L10.S1: …
L10.S20: …
Extract claims only from the target sentences above. Use the full report context for coreference resolution.
@@
Figure 7: Simplified batch extraction prompt structure and an example JSON output.
Claim Extraction Evaluation Setup.
To rigorously evaluate the claim extraction performance in Table 4, we employed an LLM-based Judge (GPT-5). Standard lexical metrics (e.g., ROUGE (Lin, 2004), Exact Match) are unsuitable for this task because the generated claims may differ in wording or granularity (e.g., one sentence split into multiple atomic claims) while preserving the same semantic meaning. We defined two key metrics: (1) Paragraph-level Semantic Recall: Measures whether each ground-truth claim is semantically covered by the extracted claims within the same source paragraph. The LLM Judge compares the ground truth claim against all candidate claims extracted from the same source sentence and determines if the core information is present. (2) Classification F1: Measures whether the LLM correctly classified the claim type for the extracted claims. The implementation code and LLM prompts used for the evaluation are included in the supplementary material.
Semantic Back-tracking for Citation Recovery
Semantic Back-tracking Evaluaton.
To validate the effectiveness of the LLM’s targeted evidence prediction, we compared it with a “Sliding Window (Patel et al., 2025)” baseline on the subset of correctly classified B/C claims ($N=131$). The sliding window method collects all citations within a window of size $k$ centered on the claim. As shown in Table 15, the proposed LLM method achieves the highest Jaccard Index (0.7070) and Precision (0.7109), outperforming the sliding window baselines ($k=5,10,15$). While increasing the window size ($k$) improves Recall (up to 0.93), it significantly degrades Precision and Jaccard due to the inclusion of irrelevant citations. This result demonstrates that the LLM’s “Evidence Position” prediction provides a precise pointer to the supporting evidence, which is crucial for efficient verification.
@@
F.3 Claim Verification
@@
In this study, the BM25 (Robertson et al., 2009) and OpenAI’s text-embedding-3-large (OpenAI, ) was used as the embedding model: The selected chunk set $K_{selected}=\{k\in K\mid rank(Sim(q,k))\leq N\}$ is combined while maintaining the original document order to form the final context $C_{final}$.
@@
The Information Verification Module’s automatic verifier is prompt-engineered to determine whether the given context supports the claim, identical to the human protocol above. The LLM infers whether the claim’s core content and numerical information match the source, then makes a final judgment.
@@
To identify the optimal configuration for the cost-efficient gpt-5-mini model, we conducted an ablation study varying Batch Size (10, 20), Reasoning Effort (Low, Medium, High), Retrieval Method (BM25 (Robertson et al., 2009), OpenAI’s text-embedding-3-large (OpenAI, )), and Context Size (Top-K=2, 4). Table 16 summarizes the results on both the Original and Adversarial (Augmented) datasets. We observed that increasing the retrieved context size from Top-K=2 to 4 improved accuracy on the Original dataset (e.g., 77.0% $\rightarrow$ 79.3% for OpenAI, Low Effort), but increased the cost by approximately 35% (\$0.95 $\rightarrow$ \$1.28). The Low Reasoning Effort setting proved to be highly cost-effective, achieving comparable or superior performance to Medium/High effort while costing significantly less. Notably, the model demonstrated high robustness on the Adversarial dataset, maintaining high accuracy (>>88%) across most configurations, suggesting that it effectively distinguishes unsupported claims even under perturbations. Consequently, prioritizing feasibility of large-scale verification (cost/throughput) while retaining strong robustness, we selected the configuration Batch 20, Low Effort, Top-K=2, and OpenAI Embedding. This setup offers a minimal cost of \$0.95 per 1k claims while maintaining a strong accuracy of 77.0% (Original) and 88.5% (Adversarial), making it the most balanced choice for our resource-constrained high-volume verification pipeline.
@@
$\Rightarrow$
Final Result: Supported
F.4 Evaluation Metrics
The evaluation metrics are designed to assess the Integrity and Sufficiency subdimensions by decomposing evidence use into complementary, claim-level signals. Rather than relying on a single aggregate score, we measure multiple failure modes of information use—factual incorrectness, unsupported attribution, unreliable or inaccessible sources, and insufficient evidence coverage—so that different weaknesses in evidence grounding can be diagnosed explicitly.
@@ F.4 Evaluation Metrics
  • Reference Diversity (Normalized HHI): Measures how evenly citations are distributed across used references using the Normalized Herfindahl-Hirschman Index (Rhoades, 1993). We define $s_{i}$ as the share of citations for reference $i$ among total citations ($s_{i}=c_{i}/\sum c)$, and HHI as follows:
    $\text{HHI}=\sum_{i=1}^{N}s_{i}^{2}$
    Based on this, the Normalized HHI score (0–10) is calculated as:
@@ 2. Aggregation
Appendix G Baseline Model Details
We use the following backbone model families in our experiments: Qwen3-235B, Gemini 2.5, Claude Opus 4.5, and GPT-5. For readability, we refer to the GPT family as GPT-5 in the paper, while the actual backbone used in our runs was GPT-5.2. Since Gemini 2.5 Pro includes reasoning by default, we use Gemini 2.5 Flash for the fast (non-reasoning) setting, and Gemini 2.5 Pro for the other settings. For think, we use each service’s default reasoning budget. For think+search, we do not build a custom retrieval pipeline; instead, we use the built-in web search system provided by each service. We exclude Gemini think+search because citation information is not provided in its outputs. For WebThinker (Li et al., 2025b), we use Qwen3-235B as an auxiliary model. For OpenAI Deep Research, we collect reports generated from the service environment as of August 2025.
Appendix H Human-Correlation Experiment Setup
We measure alignment between LLM judges and expert human judgments on 45 reports. We consider five domains and sample three tasks from each, yielding 15 tasks in total. For each task, we randomly sample three reports from those produced by five Deep Research systems—OpenAI Deep Research, Gemini 2.5 Pro Deep Research, Claude Opus 4.1 Deep Research, WebThinker, and Qwen3-235B Deep Research—resulting in $15\times 3=45$ reports overall. WebThinker (Li et al., 2025b) uses Qwen3-235B as an auxiliary model.
@@ Appendix H Human-Correlation Experiment Setup
To measure alignment with expert human judgments, we report Pearson correlation ($r$), Spearman rank correlation ($\rho$), and pairwise agreement (PA). Each task corresponds to a single query and contains three reports. For each task, we compute Pearson $r$ and Spearman $\rho$ between human and model scores across the three reports, and report the mean over the 15 tasks. Following DeepBench (Du et al., 2025), PA is the fraction of report pairs within a task for which the model’s relative preference matches the human relative preference; with three reports, each task has $\binom{3}{2}=3$ pairs. We compute PA per task and report the mean across tasks. To match the human 1–5 scale, we rescale model scores to 1–5 by dividing 1–10 scores by 2. For PA, a pair is counted as an agreement if the model and human judgments induce the same relation (>>, <<, or ==) for that pair. Tasks with undefined correlations (e.g., zero variance) are excluded from the corresponding averages.
@@
Expert contributors participated as paid contractors via a professional vendor; no sensitive personal data were collected, and participation followed the vendor’s consent and compensation policies.
Appendix I Prompts
This appendix provides the prompt templates used in our evaluation pipeline. Figure 8 shows an example evaluator prompt for the Request Fulfillment dimension, one of the report-quality dimensions. Figure 9 shows the full prompt for claim extraction and classification. Figure 9 shows the full prompt used for the claim extraction and classification task. Figure 10 shows the complete prompt for the claim verification and source reliability task.
Evaluation Prompt Template for Request Fulfillment
# 1. Overview
Evaluate the report using the provided rubric, which operationalizes the requirements of a professional, expert-level long-form
report.
For each rubric item (e.g., C1-1, C1-2, Q1-1, Q1-2, …), provide systematic evaluation reasoning (including relevant evidence
from the report) and assign an item-level score.
Scores must be integers from 1 to 10. If the report contains no assessable material for a given item, enter “N/A” instead of a numeric score.
# 2. Evaluation Method
Evaluate each rubric item strictly using the provided rubric and the Expert Evaluation Guidance (EG), and do not make arbitrary
judgments outside the rubric and EG.
# 3. Expert Evaluation Guideline (EG)
Expert Evaluation Guidance (EG) provides task-specific expert criteria: it enumerates required content elements and expert
expectations as concrete, verifiable statements that can be checked directly against the report.
The EG has absolute priority in every evaluation: each rubric item must be evaluated on the basis of the EG with all applicable EG
requirements applied in full, and if EG requirements are not met, no high score (Perfect or Excellent) may be awarded regardless of
supplementary strengths; supplementary strengths may only be considered once full compliance with the EG has been confirmed.
## Expert Evaluation Guideline
{expert guideline}
# 4. Rubric Items
[Omitted for brevity.]
# 5. How to Score: Coverage (C) vs. Quality (Q)
Each rubric item has either **C (Coverage) or Q (Quality)** attribute, and each is evaluated independently.
## 5.1 Coverage (C) Evaluation
This item is evaluated based on whether every required component is present and fully addressed wherever relevant in the report.
**Evaluation Method (Coverage/C):**
1. Identify all required elements for this rubric item.
2. For each required element, evaluate the relevant parts of the report as Pass/Fail (met/not met).
3. Classify Fails as core gaps vs minor omissions.
4. Assign a 1–10 score based on the number and type of Fails.
**Scoring Guidelines:**
[Omitted for brevity; see Table 12.]
**Core Principles:**
* Even one core gap makes Excellent (7-8) impossible
* Multiple core gaps make Good (5-6) impossible
## 5.2 Quality (Q) Evaluation
This item evaluates how well the report executes the relevant written content for the rubric item.
**Evaluation Method:**
1. Evaluate only the written parts relevant to this rubric item, and do not evaluate omissions.
2. Make an overall-level judgment (academic/professional level) based on the written content, and assign a provisional 1–10 score.
3. If a core element falls short of the provisional score level, adjust the final score downward to match the level of that core
element.
**Scoring Guidelines:**
[Omitted for brevity; see Table 12.]
**Core Principles:**
* If an EG core element falls short, the overall Q score should be lowered accordingly; weaknesses in non-core elements have limited
impact on the overall score.
* Review the quality aspects specified by the rubric item (e.g., depth, logic, rigor, precision, clarity, accuracy, balance,
methodological soundness).
# 6. Output Format
For each rubric item, write the "description" as score justification grounded in the scoring guidelines, not as a general pros/cons
summary.
(1) Band alignment: The description must match the score-band definition (Coverage or Quality) for the score you assigned.
(2) Deficiency-focused: Describe what is inadequate or insufficient with concrete specifics, grounded in the report.
(3) Consistency enforcement: Deficiency severity must be consistent with the assigned band; otherwise revise the score.
[JSON schema omitted for brevity]
[User Query]
{query}
[Expert Report]
{doc}
Figure 8: Abbreviated evaluator prompt template for Request Fulfillment. We omit the full rubric item list, detailed scoring bands, and the JSON output schema for space.
Full Prompt for Claim Extraction and Classification
You are an expert fact-checker. Your task is to extract distinct claims from the provided Target Sentences
and classify each one into one of the following categories.
Class Definitions:
| Class | Definition |
| A: Cited Claim | Claims that include explicit citations (e.g., [1] or (Kim, 2024)) directly within the sentence. |
| B: Uncited - Same Section/Paragraph | Claims whose supporting citations appear within the same section or paragraph.
Crucially, the Evidence Position MUST be in the same section or paragraph. If the evidence is in a previous section, it is Class C. |
| C: Uncited – Previous Section/Paragraph | Claims whose supporting citations appear in previous sections or paragraphs. T1
he Evidence Position MUST be in a previous section. |
| D: Uncited – Recap/Structural | Simple mentions of introduction, conclusion, abstract, or structural summaries. |
| E: Uncited – Citation not Required | Widely accepted facts, common knowledge, author’s own findings, subjective claims,
experimental results, methodological contributions, etc. WARNING: Conclusions, results, or findings of the paper are NOT Class E.
They are specific claims that require evidence (Class A/B/C) or are uncited (Class F). |
| F: No-Citation – Unknown Source | Specific factual assertions that require verification but lack any identifiable
supporting source in the text. If a claim is specific (e.g., "X has a radius of Y", "X is planar") and has no citation attached to it
or the sentence immediately preceding/following it that covers this fact, it is Class F.
Do not assume a nearby citation covers it unless it clearly does. |
Instructions:
1. Read the Report Context to understand the global context.
2. Process the "Target Sentences":
   - Break down the text into atomic claims. A single sentence may contain multiple claims
   (e.g., "X is Y, and Z requires W" -> Claim 1: "X is Y", Claim 2: "Z requires W").
   - Extract ALL statements, including facts, opinions, structural descriptions, and summaries.
3. For each extracted claim, analyze its relationship with the context and citations:
   - Step 1: Specificity Check. Contains numbers, chemical properties, specific results? -> Likely A, B, C, or F.
   - Step 2: Citation Check.
     - Citation in same sentence? -> Class A.
     - Citation in same paragraph? -> Class B.
     - Citation in previous section? -> Class C.
   - Step 3: Uncited Check.
     - Specific fact but no citation? -> Class F.
     - General knowledge/methodology/author opinion? -> Class E.
     - Structural summary? -> Class D.
4. Determine Evidence Position:
   - For Class B or C, identify the exact sentence index (e.g., "L1.S1") that contains the citation supporting this claim.
   - Must contain an explicit citation.
5. Output Format:
   - Return a JSON object with a list of claims.
   - Each claim must include: ‘position‘ (line/sent index from input), ‘claim‘ (text), ‘claim_type‘ (A-F), ‘rationale‘,
   ‘numeric‘ (bool), ‘citations‘ (list of strings), ‘implicit_citations‘ (list), ‘cross_references‘ (list).
Input Format:
# Report Excerpt
# Target Sentences
L1.S1: ...
L1.S2: ...
Extraction and Classification Logic:
- If a sentence is "Most perovskites are unstable[1], but our new material is stable.", extract TWO claims:
  1. "Most perovskites are unstable." (Class A, citations=[’1’])
  2. "Our new material is stable." (Class E - Author’s finding, or Class F if it lacks proof provided elsewhere)
Examples:
*Example Input:*
L1.S1: Several studies[1] have shown that urban green spaces can reduce ambient air temperatures by up to 2 °C. This is crucial.
*Example Output (Conceptual):*
1. Claim: "Several studies have shown that urban green spaces can reduce ambient air temperatures by up to 2 °C."
   - Class: A
   - Citations: ["1"]
   - Position: "L1.S1"
2. Claim: "This is crucial."
   - Class: B (supported by L1.S1)
   - Evidence Position: "L1.S1"
   - Position: "L1.S1"
Figure 9: Full prompt for Claim Extraction and Classification.
Full Prompt for Claim Verification
You are an expert fact-checker. Verify the following claims against the provided context.
Be extremely strict. High precision is required.
For each claim, determine if it is supported by the context.
Result should be based on semantic meaning, not just keyword matching. If the text implies the claim is true,
mark it as supported.
Verification Result (‘result‘):
- supported: The text explicitly states or clearly implies the claim is true.
- conflict: The text explicitly contradicts the claim.
- not_supported: The text does not contain enough information.
- error: Access/processing error (4xx/5xx, captcha, paywall, etc.)
[Error Cases — Details]
- Set result to "error" in the following cases:
  - HTTP 4xx/5xx errors
  - Access restrictions such as Captcha, Paywall, Authentication Required
  - Other processing error messages
- Briefly describe the specific error cause in explanation (e.g., "404 Not Found", "paywall restriction").
[Document Reliability — Top-Level Fields]
- reliable (true/false):
  - Evaluate whether the entire URL is from a trustworthy source
  - Example 1: Official statistics, academic journals, authoritative institutions -> true
  - Example 2: Personal blogs, social media posts, etc. -> false
- reliable_explanation (string or null):
  - Rationale for reliability judgment (optional)
# Examples:
Claim: "GA+ prevents degradation."
Context: "Guanidinium (GA+) ... helps passivate defects, stabilizing the structure against degradation."
Result: supported
# url: {url}
# claims:
{claims}
# Context:
{context}
Figure 10: Full prompt for Claim Verification.